If three cards are dealt from a deck of 52, what is the probability that the first card is a heart and the next two are both diamonds?
Answer by Fombitz(32388)

P(heart)=13/52=1/4
P(diamond after heart)=13/51
P(diamond after D and H)=12/50=6/25
P(H,D,D)=(1/4)(13/51)(6/25)=39/2550